OpenStudy (anonymous):

If A's income is 37.5% more than B, what percent is B less than A?

OpenStudy (amistre64):

A = B(1.375) seems to match the first part

OpenStudy (anonymous):

Why it's B(1.375) ?

OpenStudy (amistre64):

A is all of B, plus a part of B A = B + B(%) =B(1+%)

OpenStudy (anonymous):

The answer is : $$ \frac {37.5}{137.5} \times 100$$

OpenStudy (anonymous):

@amistre64:Now as I have solved it you could try my question : http://openstudy.com/updates/4ee62ad6e4b023040d4b8104 ;);)

OpenStudy (anonymous):

Why (37.5/137.5)*100 ?

OpenStudy (anonymous):

@abdul shabeer: That's for you think .. it's not really hard to proof though ;)

OpenStudy (amistre64):

A - B is the difference between them B(1.375) - B = B(1.375 - 1) = B(.375) diff/totalA = 1% of A; so *100 for total % of A \[\frac{B(.375)}{B(1.375)}*100 = A\%\] \[\frac{.375}{1.375}*100 = A\%\] \[\frac{37.5}{1.375} = A\%\]

OpenStudy (anonymous):

You can take A's income as 100 and then .. do this in elementary ways ..